OpenStudy (anonymous):

Slove for x cos(34degrees)=50/x

OpenStudy (anonymous):

\[\left\{\left\{x\to -5 \sqrt{2 \text{ Sec}\left[\frac{17 \pi }{90}\right]}\right\},\left\{x\to 5 \sqrt{2 \text{ Sec}\left[\frac{17 \pi }{90}\right]}\right\}\right\} \]